# Asymptotic expansions for Legendre functions via differential equations having coalescing turning points

### v2

@@ 5 Associated Legendre functions
where $f(x\pm i0)=\lim_{\epsilon\to 0+}f(x\pm i\epsilon)$. From Eqs. 107 and 112 we have
@@
Similarly, from Eqs. 113, 115, and 116, we have
@@
From Eqs. 119, 107, 111, 115, 116, and 117 we then find that
@@
Next, to evaluate the coefficients in the LG expansions, we first find from Eqs. 4 and 123 that
@@
both of which are consequences of $f(a,z)$ and $g(z)$ defined by Eq. 123 satisfying the conditions at the beginning of Section 2 at the singularities $z=\pm 1$ and $z=\infty$.
Now let us evaluate the coefficients given by Eqs. 21, 22, and 23. In order to do so, similarly to (52), let
@@
Then we denote $\hat{E}_{s}(a,z)=E_{s}(a,\beta)$ ($s=1,2,3,\ldots$), and so from Eqs. 21, 22, 23, 133, and 135, after some calculation, we obtain
@@
In these, we determined the proportionality constants from Eqs. 108, 115, 116, 120, 122, and 132.
Alternatively, for (143) we can find the constant as $z\to 1$ using Eqs. 114 and 130 in place of Eqs. 120 and 132 to get
@@
By comparing Eqs. 18, 29, 19, 20, and 28 with Eqs. 141, 142, 143, and 146 we can then identify
(148) $w_{n,j}(u,a,z)=\left(1-z^{2}\right)^{1/2}e^{\pm\frac{1}{2}(\mu+1)\pi i}\mathsf{Q}^{\mu}_{\nu,\pm 1}(z),$
@@
(149) $w_{n,j}(u,a,z)=\frac{\pi}{\Gamma(\nu-\mu+1)}\left(1-z^{2}\right)^{1/2}\mathsf{P}^{-\mu}_{\nu}(\pm z),$
for $j=2,4$, respectively. Furthermore, from Eqs. 141, 142, 148, and 149
@@
for $j=1,3$, respectively, and from Eqs. 143, 146, 148, and 149
@@
With this value, $\zeta=\zeta(z)$ is defined by (8), where $\xi=\xi(z)$ is given by (125) and $\alpha$ by Eq. 155. Note that when $z=x\in(-a,a)$ we have $\zeta\in(-\alpha,\alpha)$ such that
@@
on using (154).
For $z=x\in[a,1)$ we have $\zeta\in[\alpha,\infty)$, such that
@@
If we compare Eq. 152 to Eqs. 34 and 35 we infer that $\lambda\approx u-\mu-\frac{1}{2}=\nu-\mu$ as $u\to\infty$. In fact, from Eqs. 31, 119, 148, and 149 we see that $\lambda=\nu-\mu$ exactly, and hence from Eqs. 84 and 155
@@
Next, following Eqs. 81, 82, 155, and 148, define $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ by the pair of equations
@@
On employing the parabolic cylinder function Wronskian relations (43) we obtain, from solving the pair of equations Eqs. 159 and 160, our desired explicit formulas
@@
Next, in a general setting, the asymptotic expansions for $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ come from Eqs. 89 and 90. In this specific case, we can insert Eqs. 59, 60, 65, 67, 77, 79, 74, 75, 141, and 143 into Eqs. 162 and 163 (with upper signs taken), to arrive at
@@
on recalling that $\hat{\alpha}=\alpha$ (see Eq. 158), and $\mu=(1-a^{2})^{1/2}u=(1-\frac{1}{2}\alpha^{2})u$ (see Eqs. 122 and 154).
As we showed in the more general setting, by using other pairs of Eqs. 162, 163, 164, and 165, these expansions remain valid in the full domain $D$ which includes the singularities $z=\pm 1$ and $z=\pm i\infty$, and surrounds, but does not include, the turning points $z=\pm a$. The only difference are the error terms in the four quadrants, but they are all $\mathcal{O}(u^{-n})$ as $u\to\infty$ uniformly in the respective subdomains.
Next, we express Eqs. 166 and 167 in neater forms and now drop the error terms (which all have simple explicit error bounds). To this end, we have from Eqs. 168 and 169
@@
noting only odd powers of $u$ appear here. The first three terms are found to be
@@
These expansions are uniformly valid as $u\to\infty$ for $z$ lying in the principal complex plane, with points on and close to $[-a,a]$ removed. For these excluded points, we can either use the Cauchy integral method Eq. 96, or in this case a re-expansion in inverse powers of $u^{2}$, as described next.
It is readily verified that the conditions of Theorem 4.3 are met here. Thus, the following expansions hold:
(186) $A(u,a,z)\sim 1+\sum_{s=1}^{\infty}\frac{\mathrm{A}_{s}(a,z)}{u^{2s}},\quad B(u,a,z)\sim\frac{1}{u^{2}}\sum_{s=0}^{\infty}\frac{\mathrm{B}_{s}(a,z)}{u^{2s}},$
where $\mathrm{A}_{s}(a,z)$ ($s=1,2,3,\ldots$) and $\mathrm{B}_{s}(a,z)$ ($s=0,1,2,\ldots$) are analytic at $z=\pm a$, uniformly for $10\leq a\leq 1-\delta<$ 1. These coefficients can be explictly determined by formally expanding Eqs. 183 and 185 in inverse powers of $u$. For example, the first of both are given in Eqs. 188 and 189. They are also analytic at $z=0$, and using (156) we have
(187) $\zeta=\frac{a}{\alpha}z-\frac{\alpha\left(7\alpha^{2}-24\right)}{96\,a}z^{3}+\frac{\alpha^{3}\left(317\alpha^{4}-2320\alpha^{2}+4160\right)}{30720\,a^{3}}z^{5}\\ -\frac{\alpha^{5}\left(36683\alpha^{6}-413224\alpha^{4}+1537088\alpha^{2}-1881600\right)}{20643840\,a^{5}}z^{7}+\mathcal{O}\left(z^{9}\right),$
and hence from Eqs. 183 and 185
(188) $\mathrm{A}_{1}(a,z)=\frac{1}{2}\left\{\tilde{\mathcal{E}}_{1}(a,z)+d_{1}(\alpha)\right\}^{2}+\tilde{\mathcal{E}}_{2}(a,z)\\ =-\frac{1}{8\left(4-\alpha^{2}\right)^{2}}-\frac{19\alpha^{2}}{128\left(4-\alpha^{2}\right)^{3}}z^{2}-\frac{\alpha^{4}+608\alpha^{2}-304}{1536\left(4-\alpha^{2}\right)^{4}}z^{4}+\ldots,$
@@
Now, again from (156) (or (157)), we obtain for $|z-a|<1-a$
@@
The construction of the corresponding Taylor series at $z=a$ for $\mathrm{A}_{s}(a,z)$ ($s=1,2,3,\ldots$) and $\mathrm{B}_{s}(a,z)$ ($s=0,1,2,\ldots$) becomes extremely unwieldy for even a small number of terms if $a$ is kept as a general parameter. Instead, for each prescribed value of $a$, it is more practical to numerically compute the series and then discard singular terms (fractional and negative powers of $(z-a$)), since we know that they vanish, even though numerically they are not identically zero due to rounding errors.
For example, for $a=0.5$ we find numerically in Maple (with Digits set to 40) that the Taylor series for $\mathrm{A}_{1}(0.5,z)$ is given by
(192) $\frac{8.8\times 10^{-41}}{(z-0.5)^{3}}+\frac{1.4\times 10^{-40}}{(z-0.5)^{5/2}}-\frac{1.2\times 10^{-39}}{(z-0.5)^{2}}+\frac{4.8\times 10^{-40}}{(z-0.5)^{3/2}}\\ -\frac{2.2\times 10^{-39}}{z-0.1}-\frac{2.3\times 10^{-39}}{(z-0.5)^{1/2}}-0.0091223906+1.1\times 10^{-38}(z-0.5)^{1/2}\\ -0.00034392390(z-0.5)+2.7\times 10^{-38}(z-0.5)^{3/2}+0.0010293232(z-0.5)^{2}+\ldots,$
@@
(193) $\mathrm{A}_{1}(0.5,z)\approx-0.0091223906-0.00034392390(z-0.5)\\ +0.0010293232(z-0.5)^{2}+\ldots.$
@@
We summarize the main results of this section. The Legendre functions have representations Eqs. 159, 160, and 161, where $\zeta=\zeta(z)$ is defined by (8), $\xi=\xi(z)$ is given by Eqs. 125 and 126, and $\alpha$ by Eq. 155. Here, the Legendre functions $\mathsf{Q}^{\mu}_{\nu,\pm 1}(z)$ are defined by Eqs. 115 and 116, and $\mathsf{P}^{-\mu}_{\nu}(z)$ and $\mathsf{Q}^{-\mu}_{\nu}(z)$ are the analytic continuations into the complex plane of the standard Ferrers functions, and are expressible in the forms Eqs. 117 and 118.
For $z$ lying in the principal complex plane having cuts along $(-\infty,-1]$ and $[1,\infty)$, with points on and close to the interval $[-a,a]$ removed, the coefficient functions $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ possess the expansions Eqs. 182, 183, 184, and 185 as $u=\nu+\frac{1}{2}\to\infty$, uniformly for $\delta u\leq\mu\leq u$. In these, the coefficients ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are polynomials in $\beta$ and $\hat{\beta}$, as given by Eqs. 52, 53, 56, 57, 58, 72, 73, 135, 136, 137, 138, 139, 140, 170, and 171, in which $\hat{\alpha}=\alpha$. The constants $d_{2s+1}(\alpha)$ can be determined via Eqs. 168, 169, and 177 and the asymptotic expansion of the gamma function [3, Eq. 5.11.1].
For points near or in the interval $[-a,a]$ the expansions Eqs. 182, 184, and 186 are valid, where $\mathrm{A}_{s}(a,z)$ and $\mathrm{B}_{s}(a,z)$ are analytic at $z=\pm a$, and can be evaluated by expanding Eqs. 183 and 185 in inverse powers of $u$.
@@ 6 Numerical results
where $q^{-\mu}_{\nu,1}$ is the largest positive zero of $\mathsf{Q}^{-\mu}_{\nu}(x)$. In the oscillatory interval $(0,q^{-\mu}_{\nu,1})$ this positive function is close to the amplitude of $\mathsf{P}^{-\mu}_{\nu}(x)$, while coinciding with the function in the non-oscillatory interval. For example, in Fig. 3 the function $\mathsf{P}^{-\mu}_{\nu}(x)$ (solid curve) and its envelope $M(\nu,\mu,x)$ (dotted curve), together with the reflection $-M(\nu,\mu,x)$ (dash-dotted curve) and $\mathsf{Q}^{-\mu}_{\nu}(x)$ (dashed curve), are shown for $\nu=50$ ($u=101/2$) and $\mu=(1-a^{2})^{1/2}u$ ($u=\nu+\frac{1}{2}$) with $a=0.5$ and $x\in[0,1)$. In the use of (194) we found numerically for these parameter values that $q^{-\mu}_{\nu,1}=0.42542\cdots$.
@@
Next, for $z=x\in[0,1)$, on taking the first four terms in Eqs. 182, 183, 184, and 185 define the approximations
(195) $\mathcal{A}_{4}(u,a,x)=\frac{\pi^{1/4}}{u^{1/4}\sqrt{2\Gamma(u+\mu+\tfrac{1}{2})}}\left(\frac{\zeta^{2}-\alpha^{2}}{x^{2}-a^{2}}\right)^{1/4}\left\{1+\sum_{s=1}^{3}\frac{\mathrm{A}_{s}(a,x)}{u^{2s}}\right\},$
@@
(196) $\mathcal{B}_{4}(u,a,x)=\frac{\pi^{1/4}}{u^{1/4}\sqrt{2\Gamma(u+\mu+\tfrac{1}{2})}}\left(\frac{\zeta^{2}-\alpha^{2}}{x^{2}-a^{2}}\right)^{1/4}\sum_{s=0}^{3}\frac{\mathrm{B}_{s}(a,x)}{u^{2s+2}},$
where $\zeta=\zeta(x)\in[0,\infty)$ is given by Eqs. 156 and 157. Then from Eqs. 160, 195, and 196 consider the following error for $z=x\in[0,1)$ in the approximation
@@
To assess the relative accuracy of this, we computed
@@
In both cases, we used the Taylor series about $x=a$ for $|x-a|<0.08$ as described above (see Eq. 193), and computed the coefficients in Eqs. 195 and 196 directly for all other values of $x$. The plots demonstrate about 16 digits, or better, of accuracy for both values of $a$, uniformly throughout the interval.

### v3

@@ 5 Associated Legendre functions
where $f(x\pm i0)=\lim_{\epsilon\to 0+}f(x\pm i\epsilon)$. From Eqs. 117 and 122 we have
@@
Similarly, from Eqs. 123, 125, and 126, we have
@@
From Eqs. 129, 117, 121, 125, 126, and 127 we then find that
@@
Next, to evaluate the coefficients in the LG expansions, we first find from Eqs. 4 and 133 that
@@
and both of these are consequences of $f(a,z)$ and $g(z)$, as defined by Eq. 133, satisfying the conditions at the beginning of Section 2 at the singularities $z=\pm 1$ and $z=\infty$.
Now, let us evaluate the coefficients given by Eqs. 21, 22, and 23. In order to do so, similarly to (56), let
@@
Then we denote $\hat{E}_{s}(a,z)=E_{s}(a,\beta)$ ($s=1,2,3,\ldots$), and so from Eqs. 21, 22, 23, 143, and 145, after some calculation, we obtain
@@
In these, we determined the proportionality constants from Eqs. 118, 125, 126, 130, 132, and 142. Note that, similarly to the LG functions of Section 3, $W_{n,1}(u,a,z)$ and $W_{n,3}(u,a,z)$ are actually independent of $n$, but for consistency of notation, we keep this in their subscripts.
Alternatively, for (153) we can find the constant as $z\to 1$ using Eqs. 124 and 140 in place of Eqs. 130 and 142 to get
@@
On comparing Eqs. 18, 29, 19, 20, and 28 with Eqs. 151, 152, 153, and 156 we can then identify
(158) $w_{j}(u,a,z)=\left(1-z^{2}\right)^{1/2}e^{\pm\frac{1}{2}(\mu+1)\pi i}\mathsf{Q}^{\mu}_{\nu,\pm 1}(z),$
@@
(159) $w_{j}(u,a,z)=\frac{\pi}{\Gamma(\nu-\mu+1)}\left(1-z^{2}\right)^{1/2}\mathsf{P}^{-\mu}_{\nu}(\pm z),$
for $j=2,4$, respectively. Furthermore, from Eqs. 151, 152, 158, and 159
@@
for $j=1,3$, respectively, these being independent of $n$; as above for $W_{n,1}(u,a,z)$ and $W_{n,3}(u,a,z)$, we keep this in the subscripts.
Similarly, from Eqs. 153, 156, 158, and 159,
@@
With this value, $\zeta=\zeta(z)$ is defined by (8), where $\xi=\xi(z)$ is given by (135) and $\alpha$ by Eq. 165. Note that when $z=x\in(-a,a)$ ($a>0$) we have $\zeta\in(-\alpha,\alpha)$, such that
@@
on using (164). Next, $\zeta\in[\alpha,\infty)$ when $z=x\in[a,1)$, and for $a>0$ is given implicitly by
@@
Finally, $\zeta=\left\{-\ln\left(1-x^{2}\right)\right\}^{1/2}$ for $x\in(-1,1)$ when $a=0$.
If we compare Eq. 162 to Eqs. 34 and 35 we infer that $\lambda\approx u-\mu-\frac{1}{2}=\nu-\mu$ as $u\to\infty$. In fact, from Eqs. 31, 129, 158, and 159 we see that $\lambda=\nu-\mu$ exactly, and hence from Eqs. 89 and 165
@@
Next, following Eqs. 86, 87, 165, and 158, define $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ by the pair of equations
@@
On employing the parabolic cylinder function Wronskian relations (47) we obtain, from solving the pair of equations Eqs. 169 and 170, our desired explicit formulas
@@
Next, in a general setting, the asymptotic expansions for $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ come from Eqs. 94 and 95. In this specific case, we can insert Eqs. 63, 64, 70, 72, 82, 84, 79, 80, 151, and 153 into Eqs. 172 and 173 (with upper signs taken), to arrive at
@@
on recalling that $\hat{\alpha}=\alpha$ (see Eq. 168), and $\mu=(1-a^{2})^{1/2}u=(1-\frac{1}{2}\alpha^{2})u$ (see Eqs. 132 and 164).
As we showed in the more general setting, by using other pairs of Eqs. 172, 173, 174, and 175, these expansions remain valid in the full domain $D$ which includes the singularities $z=\pm 1$ and $z=\pm i\infty$, and surrounds, but does not include, the interval $[-a,a]$ containing the turning points. The only difference in the various expansions in the four quadrants are the error terms, but these are all $\mathcal{O}(u^{-n})$ as $u\to\infty$ uniformly in the respective subdomains.
Next, we express Eqs. 176 and 177 in neater forms and now drop the error terms (which all have simple explicit error bounds). To this end, we have from Eqs. 178 and 179
@@
noting that only odd powers of $u$ appear here. The first three terms are found to be
@@
These expansions are uniformly valid as $u\to\infty$ for $z$ lying in the principal complex plane having cuts along $(-\infty,-1]$ and $[1,\infty)$, with points on and close to $[-a,a]$ removed. For these excluded points, we can either use the Cauchy integral method Eq. 101, or in this case a re-expansion in inverse powers of $u^{2}$, as described next.
It can be easily verified that the conditions of Theorem 4.3 are met here, and only even inverse powers of $u$ appear in the asymptotic series. Thus, the following expansions hold
(196) $A(u,a,z)\sim 1+\sum_{s=1}^{\infty}\frac{\mathrm{A}_{2s}(a,z)}{u^{2s}},\quad B(u,a,z)\sim\frac{1}{u^{2}}\sum_{s=0}^{\infty}\frac{\mathrm{B}_{2s}(a,z)}{u^{2s}},$
where $\mathrm{A}_{2s}(a,z)$ ($s=1,2,3,\ldots$) and $\mathrm{B}_{2s}(a,z)$ ($s=0,1,2,\ldots$) are analytic at $z=\pm a$, uniformly for $10\leq a\leq 1-\delta<$ 1. These coefficients can be explictly determined by formally expanding Eqs. 193 and 195 in inverse powers of $u$. For example, the first of both are given in Eqs. 198 and 199. They are also analytic at $z=0$, and using (166) we have for $|z|<1$ and $a>0$ that
(197) $\zeta=\frac{a}{\alpha}z-\frac{\alpha\left(7\alpha^{2}-24\right)}{96\,a}z^{3}+\frac{\alpha^{3}\left(317\alpha^{4}-2320\alpha^{2}+4160\right)}{30720\,a^{3}}z^{5}\\ -\frac{\alpha^{5}\left(36683\alpha^{6}-413224\alpha^{4}+1537088\alpha^{2}-1881600\right)}{20643840\,a^{5}}z^{7}+\ldots,$
and hence from Eqs. 193 and 195
(198) $\mathrm{A}_{2}(a,z)=\frac{1}{2}\left\{\tilde{\mathcal{E}}_{1}(a,z)+d_{1}(\alpha)\right\}^{2}+\tilde{\mathcal{E}}_{2}(a,z)\\ =-\frac{1}{8\left(4-\alpha^{2}\right)^{2}}-\frac{19\alpha^{2}}{128\left(4-\alpha^{2}\right)^{3}}z^{2}-\frac{\alpha^{4}+608\alpha^{2}-304}{1536\left(4-\alpha^{2}\right)^{4}}z^{4}+\ldots,$
@@
Now, again from (166) (or (167)), we obtain for $|z-a|<1-a$ ($a>0$)
@@
Both Eqs. 197 and 200 hold in the limit as $a\to 0$ ($\alpha\to 0$), and for both one can show in this case, on referring to Eqs. 164 and 201, that $\zeta=z+\tfrac{1}{4}z^{3}+\tfrac{13}{96}z^{5}+\cdots$.
The construction of the corresponding Taylor series at $z=a$ for $\mathrm{A}_{2s}(a,z)$ ($s=1,2,3,\ldots$) and $\mathrm{B}_{2s}(a,z)$ ($s=0,1,2,\ldots$) becomes extremely unwieldy, for even a small number of terms, if $a$ is kept as a general parameter. Instead, for each prescribed value of $a$, it is more practical to numerically compute the series and then discard singular terms (fractional and negative powers of $(z-a$)), since we know that they vanish, even though numerically they are not identically zero due to rounding errors.
For example, for $a=0.5$ we find numerically in Maple (with Digits set to 40) that the Taylor series for $\mathrm{A}_{2}(0.5,z)$ is given by
(202) $\frac{8.8\times 10^{-41}}{(z-0.5)^{3}}+\frac{1.4\times 10^{-40}}{(z-0.5)^{5/2}}-\frac{1.2\times 10^{-39}}{(z-0.5)^{2}}+\frac{4.8\times 10^{-40}}{(z-0.5)^{3/2}}\\ -\frac{2.2\times 10^{-39}}{z-0.5}-\frac{2.3\times 10^{-39}}{(z-0.5)^{1/2}}-0.0091223906+1.1\times 10^{-38}(z-0.5)^{1/2}\\ -0.00034392390(z-0.5)+2.7\times 10^{-38}(z-0.5)^{3/2}+0.0010293232(z-0.5)^{2}+\ldots,$
@@
(203) $\mathrm{A}_{2}(0.5,z)\approx-0.0091223906-0.00034392390(z-0.5)\\ +0.0010293232(z-0.5)^{2}+\ldots.$
@@
We summarize the main results of this section. The Legendre functions have representations Eqs. 169, 170, and 171, where $\zeta=\zeta(z)$ is defined by (8), $\xi=\xi(z)$ is given by Eqs. 135 and 136, and $\alpha$ by Eq. 165. Here, the Legendre functions $\mathsf{Q}^{\mu}_{\nu,\pm 1}(z)$ are defined by Eqs. 125 and 126, and $\mathsf{P}^{-\mu}_{\nu}(z)$ and $\mathsf{Q}^{-\mu}_{\nu}(z)$ are the analytic continuations into the complex plane of the standard Ferrers functions, and are expressible in the forms Eqs. 127 and 128.
For $z$ lying in the principal complex plane having cuts along $(-\infty,-1]$ and $[1,\infty)$, with points on and close to the interval $[-a,a]$ removed, the coefficient functions $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ possess the expansions Eqs. 192, 193, 194, and 195 as $u=\nu+\frac{1}{2}\to\infty$, uniformly for $\delta u\leq\mu\leq u$. In these, the coefficients ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are polynomials in $\beta$ and $\hat{\beta}$, as given by Eqs. 56, 57, 60, 61, 62, 77, 78, 145, 146, 147, 148, 149, 150, 180, and 181, in which $\hat{\alpha}=\alpha$. The constants $d_{2s+1}(\alpha)$ can be determined via Eqs. 178, 179, and 187 and the asymptotic expansion of the gamma function [3, Eq. 5.11.1].
For points near or in the interval $[-a,a]$ the expansions Eqs. 192, 194, and 196 are valid, where $\mathrm{A}_{2s}(a,z)$ and $\mathrm{B}_{2s}(a,z)$ are analytic at $z=\pm a$, and can be evaluated by expanding Eqs. 193 and 195 in inverse powers of $u$.
@@ 6 Numerical results
where $q^{-\mu}_{\nu,1}$ is the largest positive zero of $\mathsf{Q}^{-\mu}_{\nu}(x)$. In the oscillatory interval $(0,q^{-\mu}_{\nu,1})$ this positive function is close to the amplitude of $\mathsf{P}^{-\mu}_{\nu}(x)$, while coinciding with the function in the non-oscillatory interval. For example, in Fig. 3 the function $\mathsf{P}^{-\mu}_{\nu}(x)$ (solid curve) and its envelope $M(\nu,\mu,x)$ (dotted curve), together with the reflection $-M(\nu,\mu,x)$ (dash-dotted curve) and $\mathsf{Q}^{-\mu}_{\nu}(x)$ (dashed curve), are shown for $\nu=50$ ($u=101/2$) and $\mu=(1-a^{2})^{1/2}u$ ($u=\nu+\frac{1}{2}$) with $a=0.5$ and $x\in[0,1)$. In the use of (204) we find numerically for these parameter values that $q^{-\mu}_{\nu,1}=0.42542\cdots$.
@@
Next, for $z=x\in[0,1)$, on taking the first four terms in Eqs. 192, 193, 194, and 195 define the approximations
(205) $\mathcal{A}_{4}(u,a,x)=\frac{\pi^{1/4}}{u^{1/4}\sqrt{2\Gamma(u+\mu+\tfrac{1}{2})}}\left(\frac{\zeta^{2}-\alpha^{2}}{x^{2}-a^{2}}\right)^{1/4}\left\{1+\sum_{s=1}^{3}\frac{\mathrm{A}_{2s}(a,x)}{u^{2s}}\right\},$
@@
(206) $\mathcal{B}_{4}(u,a,x)=\frac{\pi^{1/4}}{u^{9/4}\sqrt{2\Gamma(u+\mu+\tfrac{1}{2})}}\left(\frac{\zeta^{2}-\alpha^{2}}{x^{2}-a^{2}}\right)^{1/4}\sum_{s=0}^{3}\frac{\mathrm{B}_{2s}(a,x)}{u^{2s}},$
where $\zeta=\zeta(x)\in[0,\infty)$ is given by Eqs. 166 and 167. Then from Eqs. 170, 205, and 206 consider the following error for $z=x\in[0,1)$ in the approximation
@@
To assess the relative accuracy of this, we computed the function
@@
In both cases, we used the Taylor series about $x=a$ for $|x-a|<0.08$ as described above (see Eq. 203), and computed the coefficients in Eqs. 205 and 206 directly for all other values of $x$. The plots demonstrate about 16 digits or better of accuracy for both values of $a$, uniformly throughout the interval.